in triangle efg the measure of angle f equals 3 times the measure of angle g and the measure of angle e equals the measure of angle f minus thirty degrees find the measure of each angle
Answer by ReadingBoosters(3246)

e+f+g = 180
f=3g
e=f-30
(f-30)+3g+g = 180
((3g)-30)+3g+g=180
7g-30=180
7g=180+30
7g=210
g=30 degrees
f=3(30)=90 degrees
e=90-30=60 degrees
Proof
30+90+60 = 180